# Information Gain-Guided Causal Intervention for Autonomous Debiasing Large Language Models

### v1

@@ 4.2 Evaluation Tasks
Transfer Test Sets We evaluate the performance of LLMs in maintaining strong transferability across datasets for the same task, using SNLI (Bowman et al., 2015), MRPC (Dolan and Brockett, 2005), IMDB (Maas et al., 2011), and Squadv2 (Rajpurkar et al., 2018) as transfer test sets for NLI, PI, SA, and QA tasks, respectively.
Challenge Test Sets We assess the robustness of models against dataset bias using challenge test sets, specifically designed by manually eliminating the dataset biases from the dataset to challenge models. With the lack of corresponding challenge datasets for QA and SA task, we use HANS McCoy et al. (2019) and PAWS Zhang et al. (2019) as challenge test sets for NLI and PI task.
Additionally, to evaluate the general ability of LLMs, we utilize the extensively used MMLU (Hendrycks et al., 2021) and BBH (Suzgun et al., 2023) datasets. For BBH datasets, we use the exact match as the evaluation metric. For other datasets, we measured the accuracy to compare the results of different methods.
@@ 4.3 Baseline Methods
Llama3.1 MNLI HANS QQP PAWS
Vanilla SFT 83.5 71.3 86.6 63.4
RAZOR 83.7 72.5 86.4 62.3
IGCIDB 83.4 74.2 86.2 66.9
Gemma2 MNLI HANS QQP PAWS
Vanilla SFT 79.5 65.8 86.8 59.9
RAZOR 79.4 67.8 86.7 58.4
IGCIDB 79.1 70.4 86.5 62.6
Table 2: Comparison of IGCIDB with baselines across two LLMs on ID and challenge test sets.
Llama3.1 Gemma2
Methods MMLU BBH MMLU BBH
Vanilla SFT 60.2 42.6 68.8 44.9
RAZOR 59.3 41.7 68.9 44.3
IGCIDB 60.5 42.3 68.5 45.6
Table 3: Comparison of IGCIDB with baselines across two LLMs on the general ability benchmarks.
Prior Knowledge-Based Debiasing Method Because there are no training-based automatic debiasing methods for LLMs. So we choose the latest prior knowledge-based debiasing method RAZOR (Yang et al., 2025) and adapt it to the instruction tuning setting of LLMs. More details can be found in the Appendix E.
@@ 5.1 Generalizability on Transfer Test Sets
Task Examples Gold Baseline IGCIDB
@@
Table 4: Examples of case study. The baseline method tends to use the biased feature ‘individual’s popularity’ to answer Scottish Parliament’s member rather than an outside proposer in the first example.
(2) Compared to the vanilla supervised fine-tuning method, IGCIDB achieves consistent performance improvement on transfer test sets while maintaining performance on the in-domain datasets. This demonstrates that by leveraging the information gain-guided causal intervention, the biased features will not contribute to accurately predicting the answers for the instruction-tuning dataset. As a result, IGCIDB can effectively debias LLMs by fine-tuning on the debiased dataset to improve the generalizability of LLMs.
(3) Compared with the prior knowledge-based debiasing method RAZOR, IGCIDB shows relatively better performance on transfer test sets, which demonstrates the effectiveness of our method for debiasing LLMs. Additionally, compared to the vanilla supervised fine-tuning method, the performance of IGCIDB increases on all transfer test sets, while the performance of RAZOR decreases on the MRPC dataset. On the one hand, the performance degradation of RAZOR shows the limitations of the prior knowledge-based debiasing method that the same prior knowledge may not be applicable to all tasks. On the other hand, the performance improvement of IGCIDB on all transfer test sets further indicates the necessity of automatically identifying and eliminating dataset biases.
5.2 Generalizability on Challenge Test Sets
To investigate the generalizability of IGCIDB on the manually debiased datasets, we examine our method on HANS and PAWS datasets. As shown in Table 2, compared to the prior knowledge-based debiasing method RAZOR and the vanilla supervised fine-tuning method, IGCIDB shows better performance on both challenge test sets while maintaining performance on the in-domain datasets, which underscores the efficacy of our method for debiasing LLMs. Moreover, compared to the vanilla supervised fine-tuning method, the performance of IGCIDB increases on both challenge test sets, while the performance of RAZOR decreases on PAWS. This shows the limitations of the prior knowledge-based debiasing method that the same prior knowledge may not be applicable to all tasks, and further demonstrates the necessity of automatically identifying and eliminating dataset biases.
@@ 5.3 General Ability of LLMs
Though effectiveness for debiasing LLMs, we hope that the general ability of LLMs will not be decreased compared to the vanilla supervised fine-tuning method. Therefore, we utilize MMLU and BBH datasets to examine the general ability of different methods.
As shown in Table 3, the general ability of the IGCIDB method is not decreased and even a little better than that of the vanilla supervised fine-tuning method when using Gemma2-9B for experiments. However, the performance of RAZOR is generally lower than the vanilla supervised fine-tuning method. This further indicates the effectiveness of our method.
@@ 5.4 Analysis of Model Confidence
To further investigate the effectiveness of the IGCIDB method, we use the kernel density estimation to draw the distributions of model confidence on bias-conflicting samples. Specifically, we conduct experiments on the HANS dataset, where there is a high lexical overlap rate between the premise and hypothesis across all data. McCoy et al. (2019) have demonstrated that for the NLI task, LLMs tend to predict entailment rather than contradiction when there is a high lexical overlap rate between the premise and hypothesis. Therefore, we consider the samples with label contradiction as the bias-conflicting samples.
@@ 5.5 Case Study
To demonstrate the efficacy of the proposed method, we present a case study for QA task in Table 4 (more case studies can be seen in the Appendix F). We compare our proposed method based on Llama3.1-8B with the baseline of the vanilla supervised fine-tuning method.
From Table 4, we can find that the baseline method tends to use the biased feature “individual’s popularity” to answer Scottish Parliament’s member in both cases. However, since the given context explicitly states “a private bill can be submitted to Parliament by an outside proposer”, the correct answer is “an outside proposer” which is less popular than “Scottish Parliament’s member” in the first example. Combining the case studies in the Appendix F, it can be concluded that our method IGCIDB, exhibits enhanced robustness compared to the baseline approach. The resilience of our method to conditions with different biased features suggests that it is well-suited for real-world environments where reliable LLMs are crucial.
@@ 6 Related Work
LLMs Debiasing Previous works demonstrate that LLMs still suffer from biases such as negation bias Mukherjee et al. (2021) and popularity bias Klimashevskaia et al. (2024). To debias LLMs, a series of methods depend on researchers’ prior knowledge to artificially detect the potential dataset biases, followed by debiasing through prompt-based regularization, instruction tuning, or preference optimization Wang et al. (2023); Ganguli et al. (2023); Tang et al. (2024). However, these methods are limited by the reliance on researchers’ prior knowledge. Additionally, due to the diversity of dataset biases Poliak et al. (2018); Schuster et al. (2019); Schick et al. (2021), it is impractical to identify them one by one manually for all tasks and datasets.
To address these issues, previous works proposed automatic debiasing methods. They train biased models to automatically extract biased features which characterizes the dataset biases Utama et al. (2020b); Du et al. (2023); Sanh et al. (2020); Lyu et al. (2023), and then some methods such as confidence regularization (Utama et al., 2020a) are utilized for regularizing the main model. However, such methods are designed for discriminative models and are difficult to adapt to generative LLMs. To automatically debias generative LLMs, Sun et al. (2024) proposed the causal-guided active learning method, which first automatically induces explainable biased features, and then debias LLM utilizing in-context learning. However, this method is based on in-context learning so that it can only suppress bias in LLMs but cannot completely eliminate it in theory. Moreover, due to the limited instruction following ability of LLMs, it can only address limited numbers of biases. For example, it can only suppress two biases when using Llama2-13b-Instruct.
@@
Causal Mechanism A series of recent works have explored to debias the neural network models based on the causal mechanism. Lv et al. (2022) utilized Structural Causal Model (SCM) to model auxiliary variables and exhibit promising performance in domain generalization. Sun et al. (2024) explored the causal invariance mechanism to automatically identify biased samples and biased features. In this work, we use the causal intervention (Pearl, 2009) to automatically debias LLMs.
@@ 7 Conclusions
In this paper, we propose an information gain-guided causal intervention debiasing framework. According to the condition of unbiased dataset that the biased feature should not provide any additional information for predicting the answers, we can derive the data debiasing goal using information gain theory. Then we utilize a causal intervention-based data rewriting method for achieving the data debiasing goal. Finally, a standard supervised fine-tuning process is used to debias LLMs. Experimental results show that our approach can effectively debias LLMs to enhance their generalizability in both transfer and challenge settings.
8 Limitations
@@
Appendix A Proof of the information Gain-guided Data Debiasing Goal
From Eq. 1, we can know that:
$H(Y|B)=H(Y),$ (5)
where $H$ is the information entropy. The entropy $H(Y)$ and the conditional entropy $H(Y|B)$ over all possible values of $Y$ and $B$ are defined as:
@@ Appendix A Proof of the information Gain-guided Data Debiasing Goal
Because $p(y|b)$ is always greater than or equal to $P(y)$, In order for Eq. 9 to hold, it is necessary to ensure that $P(y)=P(y|b)$ for any y and b, so that we have
$P(Y)=P(Y|B),$ (11)
@@
Appendix B Data Debiasing Goal Based on Different Biased Features
For each task, we choose the most general biased features identified by the CAL method for data debiasing. Specifically, for NLI and PI task, the identified biased feature is the lexical overlap between the two sentences. However, the lexical overlap rate can be any value between 0 and 1. To establish the data debiasing goal, we manually defined the value space of this biased feature: low lexical overelap (for lexical overelap rate which is less than 0.4), medium lexical overelap (between 0.4 and 0.6), low lexical overelap (bigger than 0.6). For these two tasks, the value space of the answer is easy to enumerate, so no modifications are necessary. For the SA task, the identified bias feature is whether the negation word exists in the sentence, so the data debiasing goal does not need to be adjusted as illustrated in Sec. 3.2. For the bias identification in the QA task, we choose two primary categories of questions (as shown in D): those aim at inquiring about a person and those ask for a location. From these two categories, the identified biased features are the popularity of individual and the popularity of location respectively. For the biased feature popularity of the individual, the adjustment process has been shown in Sec. 3.2. For the biased feature popularity of the location, the adjustment process is similar.
Task Examples Gold Baseline IGCIDB
@@ Appendix B Data Debiasing Goal Based on Different Biased Features
Appendix C Details of the biased feature checking process
To ensure that the biased feature of the rewritten samples satisfies the rewriting objectives, we employ different metrics to check these biased features. For lexical overlap bias, we calculate the lexical overlap rate between two input sentences. For negation word bias, we directly verify the presence of any predefined negation words in the inputs. For the popularity bias, we utilize the biased LLM to check if the individual or the location is considered popular.
Appendix D More Experimental Details
When applying the IGCIDB method to QA tasks, we focus on two primary categories of questions for data debiasing due to the CAL method’s requirement for uniformity in the question’s objective. One category involves questions aimed at inquiring about a person, while the other involves questions about a location. After data debiasing, we also integrate those unchanged data (such as the Lima dataset) into the debiased dataset. For the evaluation of the MNLI, QQP, MRPC, SST2, Squadv1, and Squadv2 datasets, we utilize the validation sets for evaluation due to the unavailability of the test sets. For the other datasets, we use the test sets for evaluation.

### v2

@@ 4.2 Evaluation Tasks
Transfer Test Sets We evaluate the performance of LLMs in maintaining strong transferability across datasets for the same task, using SNLI [2], MRPC [6], IMDB [19], and Squadv2 [28] as transfer test sets for NLI, PI, SA, and QA tasks, respectively.
Challenge Test Sets We assess the robustness of models against dataset bias using challenge test sets, specifically designed by manually eliminating the dataset biases from the dataset to challenge models. With the lack of corresponding challenge datasets for QA and SA task, we use HANS [20] and PAWS [48] as challenge test sets for NLI and PI task.
Furthermore, to evaluate the general ability of LLMs, we utilize the extensively used MMLU [10], BBH [36], and TruthfulQA [15] datasets. For BBH datasets, we use the exact match as the evaluation metric. For other datasets, we measured the accuracy to compare the results of different methods.
@@ 4.3 Baseline Methods
Table 1: Comparison of ICD with baselines across two LLMs on ID and transfer test sets.
NLI PI SA QA
Llama MNLI SNLI QQP MRPC SST2 IMDB Squadv1 TriviaQA Squadv2
Vanilla 84.3$\pm$0.2 71.2$\pm$0.6 86.3$\pm$0.1 75.3$\pm$0.4 95.4$\pm$0.2 95.4$\pm$0.1 87.8$\pm$0.1 65.6$\pm$0.1 86.6$\pm$0.1
Razor 84.5$\pm$0.2 72.4$\pm$0.5 85.5$\pm$0.1 75.0$\pm$0.5 95.3$\pm$0.1 95.5$\pm$0.1 87.8$\pm$0.1 65.6$\pm$0.1 86.7$\pm$0.1
Ours 83.8$\pm$0.3 75.2$\pm$0.6 85.7$\pm$0.1 77.4$\pm$0.2 95.4$\pm$0.1 95.5$\pm$0.1 87.9$\pm$0.1 65.8$\pm$0.1 87.0$\pm$0.1
Gemma MNLI SNLI QQP MRPC SST2 IMDB Squadv1 TriviaQA Squadv2
Vanilla 76.2$\pm$0.2 63.6$\pm$0.6 87.0$\pm$0.1 73.4$\pm$0.4 96.2$\pm$0.2 95.4$\pm$0.1 85.8$\pm$0.2 68.9$\pm$0.1 84.9$\pm$0.1
Razor 76.3$\pm$0.4 64.3$\pm$0.5 86.5$\pm$0.1 72.5$\pm$0.4 96.1$\pm$0.2 95.5$\pm$0.1 85.6$\pm$0.2 68.9$\pm$0.2 85.0$\pm$0.1
Ours 75.7$\pm$0.3 65.8$\pm$0.6 86.5$\pm$0.1 75.4$\pm$0.5 96.2$\pm$0.2 95.7$\pm$0.1 85.5$\pm$0.1 69.0$\pm$0.2 85.2$\pm$0.2
@@
Table 2: Comparison of ICD with baselines across two LLMs on ID and challenge test sets.
@@
Llama MNLI HANS QQP PAWS
Vanilla 84.3$\pm$0.2 69.4$\pm$0.2 86.3$\pm$0.1 62.2$\pm$0.2
Razor 84.5$\pm$0.2 70.1$\pm$0.3 85.5$\pm$0.1 61.9$\pm$0.2
Ours 83.8$\pm$0.3 72.3$\pm$0.4 85.7$\pm$0.1 65.3$\pm$0.2
Gemma MNLI HANS QQP PAWS
Vanilla 76.2$\pm$0.2 65.5$\pm$0.3 87.0$\pm$0.1 57.1$\pm$0.2
Razor 76.3$\pm$0.4 65.7$\pm$0.3 86.5$\pm$0.1 56.6$\pm$0.3
Ours 75.7$\pm$0.3 67.0$\pm$0.3 86.5$\pm$0.1 60.5$\pm$0.4
Table 3: Comparison of ICD with baselines across two LLMs on the general ability benchmarks.
Llama MMLU BBH TruthfulQA
Vanilla 60.3$\pm$0.3 41.5$\pm$0.4 44.7$\pm$0.3
Razor 59.4$\pm$0.3 40.6$\pm$0.3 45.6$\pm$0.4
Ours 61.0$\pm$0.3 42.3$\pm$0.3 52.8$\pm$0.2
Gemma MMLU BBH TruthfulQA
Vanilla 68.3$\pm$0.1 45.3$\pm$0.1 24.1$\pm$0.2
Razor 68.4$\pm$0.2 44.9$\pm$0.2 23.5$\pm$0.3
Ours 68.8$\pm$0.2 46.9$\pm$0.3 27.3$\pm$0.3
Prior Knowledge-Based Debiasing Method Because there are no training-based automatic debiasing methods for LLMs. So we choose the latest prior knowledge-based debiasing method Razor [47] and adapt it to the instruction tuning setting of LLMs. More details can be found in the Appendix E.
@@ 5.1 Generalizability on Transfer Test Sets
(2) Compared to the vanilla supervised fine-tuning method, ICD achieves consistent performance improvement on transfer test sets while maintaining performance on the in-domain datasets. This demonstrates that by leveraging the information gain-guided causal intervention, the biased features will not contribute to accurately predicting the answers for the instruction-tuning dataset. As a result, ICD can effectively debias LLMs by fine-tuning on the debiased dataset to improve the generalizability of LLMs.
(3) Compared with the prior knowledge-based debiasing method Razor, ICD shows relatively better performance on transfer test sets, which demonstrates the effectiveness of our method for debiasing LLMs. Additionally, compared to the vanilla supervised fine-tuning method, the performance of ICD increases on all transfer test sets, while the performance of Razor decreases on the MRPC dataset. On the one hand, the performance degradation of Razor shows the limitations of the prior knowledge-based debiasing method that the same prior knowledge may not be applicable to all tasks. On the other hand, the performance improvement of ICD on all transfer test sets further indicates the necessity of automatically identifying and eliminating dataset biases.
5.2 Generalizability on Challenge Test Sets
To investigate the generalizability of ICD on the manually debiased datasets, we examine our method on HANS and PAWS datasets. As shown in Table 3, compared to the prior knowledge-based debiasing method Razor and the vanilla supervised fine-tuning method, ICD shows better performance on both challenge test sets while maintaining performance on the in-domain datasets, which underscores the efficacy of our method for debiasing LLMs. Moreover, compared to the vanilla supervised fine-tuning method, the performance of ICD increases on both challenge test sets, while the performance of Razor decreases on PAWS. This shows the limitations of the prior knowledge-based debiasing method that the same prior knowledge may not be applicable to all tasks, and further demonstrates the necessity of automatically identifying and eliminating dataset biases.
Table 4: Examples of case study. The baseline method tends to use the biased feature ‘individual’s popularity’ and to answer Scottish Parliament’s member rather than an outside proposer in the first example. And it also utilizes the biased feature ‘high lexical overlap’ to answer that the two sentences have thesame meaning in the third example.
Task Examples Gold Baseline ICD
@@ 5.3 General Ability of LLMs
Though effectiveness for debiasing LLMs, we hope that the general ability of LLMs will not be decreased compared to the vanilla supervised fine-tuning method. Therefore, we utilize MMLU, BBH, and TruthfulQA datasets to examine the general ability of different methods.
As shown in Table 3, the general ability of the ICD method is not decreased and even better than that of the vanilla supervised fine-tuning method in general, especially on TruthfulQA and BBH datasets. However, the performance of Razor is generally lower than the vanilla supervised fine-tuning method. This further indicates the effectiveness of our method.
@@ 5.4 Analysis of Model Confidence
To further investigate the effectiveness of the ICD method, we use the kernel density estimation to draw the distributions of model confidence on bias-conflicting samples. Specifically, we conduct experiments on the HANS dataset, where there is a high lexical overlap rate between the premise and hypothesis across all data. McCoy et al. [20] has demonstrated that for the NLI task, LLMs tend to predict entailment rather than contradiction when there is a high lexical overlap rate between the premise and hypothesis. Therefore, we consider the samples with label contradiction as the bias-conflicting samples.
@@ 5.5 Case Study
To demonstrate the efficacy of the proposed method, we present a case study for QA ans PI task in Table 4 (more case studies can be seen in the Appendix F). We compare our proposed method based on Llama3.1-8B with the baseline of the vanilla supervised fine-tuning method.
From Table 4, we can find that the baseline method tends to use the biased feature “individual’s popularity” to answer Scottish Parliament’s member in both cases of the QA task. However, since the given context explicitly states “a private bill can be submitted to Parliament by an outside proposer”, the correct answer is “an outside proposer” which is less popular than “Scottish Parliament’s member” in the first example. Similarly, in the context of the PI task, the baseline method also tends to utilize the biased feature “high lexical overlap” to answer the meaning of the two sentences to be the same, failing to comparing the meaning of the two sentences.
However, after applying the ICD method, our approach is capable of avoiding making inferences based on these biased features, thereby enabling the LLM to arrive at the correct answers. Combining the case studies in Appendix F, it can be concluded that our method ICD, exhibits enhanced robustness compared to the baseline approach. The resilience of our method to conditions with different biased features suggests that it is well-suited for real-world environments where reliable LLMs are crucial.
@@ 6 Related Work
LLMs Debiasing Previous works demonstrate that LLMs still suffer from biases such as negation bias [21] and popularity bias [13]. To debias LLMs, a series of methods depend on researchers’ prior knowledge to artificially detect the potential dataset biases, followed by debiasing through prompt-based regularization, instruction tuning, or preference optimization [42, 9, 37]. However, these methods are limited by the reliance on researchers’ prior knowledge. Additionally, due to the diversity of dataset biases [26, 31, 30], it is impractical to identify them one by one manually for all tasks and datasets.
To address these issues, previous works proposed automatic debiasing methods. They train biased models to automatically extract biased features which characterizes the dataset biases [41, 7, 29, 18], and then some methods such as confidence regularization [40] are utilized for regularizing the main model. However, such methods are designed for discriminative models and are difficult to adapt to generative LLMs. To automatically debias generative LLMs, Sun et al. [35] proposed the causal-guided active learning method, which first combines the active learning and causal mechanisms to automatically induces explainable biased features, and then debias LLM utilizing in-context learning. However, this method is based on in-context learning so that it can only suppress bias in LLMs but cannot completely eliminate it in theory. Moreover, due to the limited instruction following ability of LLMs, it can only address limited numbers of biases. For example, it can only suppress two biases when utilizing Llama2-13b-Instruct [39].
@@
Causal Mechanism Recently, a series of recent works have explored to debias the neural network models based on the causal mechanism. Lv et al. [17] utilized Structural Causal Model (SCM) to model auxiliary variables and exhibit promising performance in domain generalization. Xu et al. [45] uses the causal effect estimation method to debias the language models for the fact verification task. Lim et al. [14] applies the causal mediation analysis for debiasing NLU models. Sun et al. [35] explored the causal invariance mechanism to automatically identify biased samples and biased features. In this work, we combine information theory and the causal intervention [25] method to automatically debias LLMs.
@@ 7 Conclusions
In this paper, we propose an information gain-guided causal intervention debiasing framework. According to the condition of an unbiased dataset that the biased feature should not provide any additional information for predicting the answers, we can derive the data debiasing goal using information theory. Then we utilize a causal intervention-based data rewriting method for achieving the data debiasing goal. Finally, a standard supervised fine-tuning process is used to debias LLMs. Experimental results show that our approach can effectively debias LLMs to enhance their generalizability in both transfer and challenge settings.
Appendix A Limitations
@@
Appendix B Proof of the information Gain-guided Data Debiasing Goal
The entropy $H(Y)$ and the conditional entropy $H(Y|B)$ over all possible values of $Y$ and $B$ are defined as:
@@ Appendix B Proof of the information Gain-guided Data Debiasing Goal
Appendix C Data Debiasing Goal Based on Different Biased Features
For each task, we choose the most prominent biased features (the first biased feature listed in Appendix. G) identified by the CAL method for data debiasing. Specifically, for NLI and PI tasks, the most prominent biased feature is the lexical overlap between the two sentences. However, the lexical overlap rate can be any value between 0 and 1. To establish the data debiasing goal, the value space of this biased feature is divided into 3 classes: low lexical overelap (for lexical overelap rate which is less than 0.4), medium lexical overelap (between 0.4 and 0.6), low lexical overelap (bigger than 0.6). For these two tasks, the value space of the answer is easy to enumerate, so no modifications are necessary. For the SA task, the most prominent biased feature is whether the negation word exists in the sentence, so the data debiasing goal does not need to be regularized as illustrated in Sec. 3.2. For the QA task, the most prominent biased feature is the popularity of a person or a location, and the regularization process has been shown in Sec. 3.2.
Task Examples Gold Baseline ICD
@@ Appendix C Data Debiasing Goal Based on Different Biased Features
Appendix D Details of the biased feature checking process
To ensure that the biased feature of the rewritten samples satisfies the rewriting objectives, we employ different metrics to check these biased features. For lexical overlap bias, we calculate the lexical overlap rate between two input sentences. For negation word bias, we directly verify the presence of any predefined negation words in the inputs. For the popularity bias, we utilize the biased LLM to check if the individual or location is considered popular.